OpenStudy (camerondoherty):

9. ∠ABC and ∠EBF are a pair of vertical angles; m∠ABC = 3x + 8 and m∠EBF = 2x + 48. What are m∠ABC and m∠EBF?

OpenStudy (shadowlegendx):

Answer choices? And what do you think the answer is?

OpenStudy (camerondoherty):

No answer choices... But i think youre supposed to solve for x: 3x + 8 = 2x + 48 Then apply x to each equation and find the measures...

OpenStudy (shadowlegendx):

Do you know how to solve for x, if this is the path we are supposed to go down?

OpenStudy (camerondoherty):

Yep ik 3x + 8 = 2x + 48 -2x -2x x + 8 = 48 x = 40

OpenStudy (shadowlegendx):

Now what would you do with x = 40 ???

OpenStudy (camerondoherty):

m∠ABC = 3x + 8 3(40) + 8 120 + 8 m∠ABC = 128 m∠EBF = 2x + 48 2(40) + 48 80 + 48 128

OpenStudy (shadowlegendx):

Ugh I keep getting disconnected

OpenStudy (camerondoherty):

Lol me too

OpenStudy (shadowlegendx):

It's like a roller coaster

OpenStudy (shadowlegendx):

Ok so now you got the answer for both angles ;3

OpenStudy (camerondoherty):

I did it through the right method?

OpenStudy (shadowlegendx):

Perhaps c; I had 20min of sleep so I may not be the best source for mathematical help xD

OpenStudy (camerondoherty):

XD

OpenStudy (shadowlegendx):

But I do agree with how you solved this

OpenStudy (shadowlegendx):

Very well done

OpenStudy (camerondoherty):

Thanks :3